OpenStudy (anonymous):

00:51:14 Use the distributive property to simplify the expression. –7(8x – 3) + x

OpenStudy (anonymous):

(x-7 (-3+8 x)) dx = 21 x-(55 x^2)/2+constant x = 21/55.

OpenStudy (anonymous):

-56x+21+x -55x+21